1. What is the future value in 7.4 years of $8651 invested in an account with a stated annual interest rate of 10%, compounded annually?

2. What is the future value in 14 years of $818 invested in an account with a stated annual interest rate of 4.3%, compounded semi-annually?

3. What is the future value in 5 years of $7922 invested in an account with a stated annual interest rate of 5.8%, compounded monthly

4. What is the future value in 13 years of $8081 invested in an account with a stated annual interest rate of 3.9%, compounded continuously?
Answer #1

FV = PV(1+r/t)^nt

1. FV = 8651(1+0.1)^7.4 = 17513.46917 i.e $17513.47

2. FV = 818(1+0.043/2)^(14*2) = 818(1+0.0215)^28 = $1483.97878 i.e 1483.98

3. FV = 7922(1+0.058/12)^(5*12) = 7922(1+0.004833)^60 = $10579.7851 i.e $10579.79

4. FV in case of continuous compounding = Pe^rt = 8081*e^(0.039*13) = 8081*e^0.507 = 8081 * 1.6603028 = $13416.906988 i.e. $13416.91